Calculate the pH of a solution in which exactly 55 ml of 0.1320 M Mg(OH)2 (aq) diluted to a final volume of 287 mL. Enter the pH with 2 decimal places.
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 0.1320 M

V1 = 55 mL

V2 = 287 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (0.132*55)/287

M2 = 0.0253 M

This is final concentration of Mg(OH)2

Use:

[OH-] = 2*[Mg(OH)2]

=2*0.0253 M

= 0.0506 M

use:

pOH = -log [OH-]

= -log (5.06*10^-2)

= 1.2958

use:

PH = 14 - pOH

= 14 - 1.2958

= 12.7042

Answer: 12.70